DiRef, A DATABASE OF REFERENCES ASSOCIATED WITH THE SPECTRA OF DIATOMIC MOLECULES

\underline{P. F. BERNATH} and S. McLEOD, Department of Chemistry, University of Waterloo, Waterloo, Ontario, CANADA N2L 3G1.

For some years now we have been compiling references of diatomic molecules with the ultimate (but still distant) goal of a revision of Huber and Herzberg’s classic book ”Constants of Diatomic Molecules.” Our work is based largely on the Berkeley Newsletters, but with additions and corrections. DiRef (Diatomic Reference Database) currently consists of about 30,000 references including titles covering the period of 1974 - 2000. The starting year was chosen to ensure adequate overlap with the references found in Huber and Herzberg. An on-going project (about half-completed) is the addition of all references in Huber and Herzberg to DiRef. We endeavor to keep DiRef current through regular up-dates several times per year. DiRef is available on the Web at http://diref.uwaterloo.ca/.